OpenStudy (anonymous):

the pair of points is on the graph of an inverse variation. find the missing value (2, 8) and (x, 3)

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

@dumbcow

OpenStudy (dumbcow):

\[y = \frac{k}{x}\] plugging in point (2,8) you get k = 16 \[3 = \frac{16}{x} \rightarrow x = \frac{16}{3}\]